# Classification of thin Jordan schemes

### v2

@@ 1. Introduction
It is well known that all finite groups form a special subclass of association schemes, called thin schemes [19]. Such schemes may be characterized as those that have a maximum rank-to-order ratio among all association schemes. In this paper, we classify all Jordan schemes that attain maximum rank-to-order ratio. We further show that regular Jordan schemes with this property are in a one-to-one correspondence with a special class of Moufang loops.
The paper is organized as follows. Section 2 contains all necessary definitions and basic facts about coherent Jordan configurations and Jordan schemes. Section 3 is the most central part of our paper. There we show that the rank-to-order ratio of a Jordan scheme is bounded by $3/2$ with equality occurring if and only if the scheme is thin and nonregular. All such Jordan schemes are classified in Subsection 3.1. Subsection 3.2 treats thin regular Jordan schemes. There we establish a one-to-one correspondence between such schemes and ring-alternative Moufang loops (RA-loops). Finally, in Section 4 we show that Jordan schemes corresponding to RA-loops are autonomous, that is to say, they are not algebraic fusions of coherent configurations. Thus we describe the first infinite family of autonomous Jordan schemes. Previously, the only known autonomous Jordan scheme was the one on $15$ points described in [14] (see also [11, Section 4]) related to one of the two Siamese $\mathrm{STS}(15)$ constructed in [13, Section 4]. The fact that this Jordan scheme is autonomous was established using the COCO2P package [12] in GAP [6] and Ziv-Av’s catalog of small coherent configurations [20].
@@ 2.1. Binary relations and matrices
Let $\Omega$ be a nonempty finite set. A binary relation on $\Omega$ is an arbitrary subset $s\subseteq\Omega\times\Omega$. We refer to the set $s{\omega}:=\{\beta\in\Omega\mid(\omega,\beta)\in s\}$ as the $s$-neighborhood of $\omega\in\Omega$. A relation is called regular if $|s\omega|$ does not depend on $\omega$. In such case, we refer to the constant $n_{s}:=|s\omega|$ as the valency of $s$. If $s$ is a regular relation of valency one, then $s{\omega}$ contains a unique element which we denote as $s(\omega)$, i.e. $s{\omega}=\{s(\omega)\}$. Thus $s$ is a function on $\Omega$. All functions in this paper, including permutations from ${\sf Sym}(\Omega)$, will be considered as binary relations.
Every relation $s$ admits a transposed relation $s^{t}=\{(\alpha,\beta)\in\Omega^{2}\mid(\beta,\alpha)\in s\}$. Throughout, we will denote the unary operator that transposes relations by $\tau$, that is, $\tau:s\mapsto s^{t}$. Following [5] we call a relation $s\subseteq\Omega\times\Omega$ thin if $|s\omega|\leq 1\geq|s^{t}\omega|$ holds for each $\omega\in\Omega$.
@@
The algebra of all $\Omega\times\Omega$ matrices over a field $\mathbb{F}$ will be denoted as $M_{\Omega}(\mathbb{F})$. It is always assumed that ${\mathsf{c}har}({\mathbb{F}})\neq 2$. The standard product of two matrices $A,B\in M_{\Omega}(\mathbb{F})$ will be written as $A\cdot B$ or $AB$. while their Jordan product $\frac{1}{2}(AB+BA)$ will be denoted by $A\star B$. The Schur-Hadamard (or entry-wise) product will be denoted as $A\circ B$. The identity matrix and all-ones matrix will be denoted as $I_{\Omega}$ and $J_{\Omega}$, respectively. The linear span of any set of matrices ${\mathcal{S}}\subseteq M_{\Omega}({\mathbb{F}})$ will be denoted by $\mathbb{F}[\mathcal{S}]$. In the case where $\mathcal{S}$ is given as a family $\{A_{i}\}_{i\in I}$, we also write $\langle A_{i}\rangle_{i\in I}$ for $\mathbb{F}[\mathcal{S}]$.
@@ 2.2. Rainbow algebras
Following [8] we call a pair $\mathfrak{S}=(\Omega,S)$ a rainbow if it satisfies these conditions. The elements of $S$ are called basic relations of the rainbow. The cardinalities $|\Omega|$ and $|S|$ are called the order and rank of the rainbow, respectively. Observe that a rainbow has rank one if and only if $|\Omega|=1$. For $|\Omega|\geq 2$ one has $|S|\geq 2$, where equality is attained if and only if $S=\{1_{\Omega},\,\Omega^{2}\setminus 1_{\Omega}\}$. Such rainbows are called trivial. The adjacency matrix of a rainbow ${\mathfrak{S}}=(\Omega,S)$ is the $\Omega\times\Omega$ matrix $A({\mathfrak{S}})$ defined via $A({\mathfrak{S}})_{\alpha,\beta}=S(\alpha,\beta)$ where $S(\alpha,\beta)$ is the unique basic relation containing $(\alpha,\beta)\in\Omega^{2}$.
@@ 2.3. Coherent Jordan algebras
Recall that a homogeneous coherent configuration is called an association scheme (see [19]). Following this pattern, we will call a homogeneous coherent J-configuration a Jordan scheme. While it is well known that association schemes are always regular, Jordan schemes do not possess this property in general. It was shown in [14] that a symmetric Jordan scheme is always regular. Regarding nonsymmetric Jordan schemes, they are not necessarily regular, but have a very special structure — see [14, Proposition 2.2]. Examples of nonregular Jordan schemes also appear in Subsection 3.1 of this paper.
@@ 2.5. Algebraic fusions of J-configurations
    The relations $\Phi(s),s\in S$ form a Jordan configuration on $\Omega$.
@@
Proof of (a). One readily checks that for any pair $r,s\in S$,
$s=r^{t}\iff(\underline{s}\star\underline{r})\circ I_{\Omega}\neq 0.$
@@
$(\underline{s}\star\underline{r})\circ I_{\Omega}\neq 0\iff(\underline{\phi(s)}\star\underline{\phi(r)})\circ I_{\Omega}\neq 0,$
@@
Proof of (b). It follows from part (a) that
@@
Proof of (c). Denote ${\mathcal{A}}:=\mathbb{C}[S]$ and extend each $\phi\in\Phi$ on $\mathcal{A}$} linearly via $\phi(\underline{s}):=\underline{\phi(s)}$. We will use the same notation for the induced linear operator on $\mathcal{A}$}. Then
@@ 2.6. Constructing Jordan configurations from coherent ones
If ${\mathfrak{S}}=(\Omega,S)$ is a coherent configuration, then it is also a Jordan configuration. Recall that a permutation $\phi$ of $S$ is called an algebraic automorphism (anti-automorphism) of $\mathfrak{S}$} if its linear extension to $\mathcal{A}:=\langle\underline{s}\rangle_{s\in S}$ is an automorphism (anti-automorphism) of $(\mathcal{A},\cdot)$. The group $\mathsf{AAut}({\mathfrak{S}})$ of all algebraic automorphisms of $\mathfrak{S}$} is contained in $\mathsf{JAut}({\mathfrak{S}})$. Also, $\tau\in\mathsf{JAut}({\mathfrak{S}})$. Every anti-automorphism of $\mathfrak{S}$} is a composition of $\tau$ and some $\phi\in\mathsf{AAut}({\mathfrak{S}})$. All automorphisms and anti-automorphisms of $\mathfrak{S}$} form a subgroup of $\mathsf{JAut}({\mathfrak{S}})$ which will be denoted by $\mathsf{TAut}({\mathfrak{S}})$. If $\mathfrak{S}$} is commutative, then $\mathsf{AAut}({\mathfrak{S}})=\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$. If however $\mathfrak{S}$} is not commutative, then $\tau\not\in\mathsf{AAut}({\mathfrak{S}})$ and $\mathsf{TAut}({\mathfrak{S}}):\mathsf{AAut}({\mathfrak{S}})]=2$. It is easy to check that $\mathsf{TAut}({\mathfrak{S}})=\mathsf{AAut}({\mathfrak{S}})\times\langle\tau\rangle$. In general, the inclusion $\mathsf{TAut}({\mathfrak{S}})\leq\mathsf{JAut}({\mathfrak{S}})$ may be proper. The smallest examples are $\mathrm{CC}(8,31)$ and $\mathrm{CC}(8,32)$ of order $8$. (Enumeration of coherent configurations here is taken from Ziv-Av’s catalog [20] of small coherent configurations.
If ${\mathfrak{S}}=(\Omega,S)$ is a thin association scheme, then $S$ is a regular subgroup of ${\sf Sym}(\Omega)$. In this case a permutation $\phi\in{\sf Sym}(S)$ belongs to $\mathsf{JAut}({\mathfrak{S}})$ if and only if it satisfies $\phi(st)\in\{\phi(s)\phi(t),\phi(t)\phi(s)\}$ for all $s,t\in S$. It was shown in [15] that any such map is either an automorphism or anti-automorphism111 The authors are grateful to R. Lipyanski who informed them of the paper [15].. Thus for thin association schemes one always has $\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$.
@@ 3. Thin Jordan schemes
By [14, Proposition 2.2] there exists a partition $\Omega=\Omega_{1}\cup\Omega_{0}$ with $|\Omega_{1}|=|\Omega_{0}|$ such that every regular relation is contained in $\Omega_{0}^{2}\cup\Omega_{1}^{2}$ and each nonregular one is contained in either $\Omega_{0}\times\Omega_{1}$ or $\Omega_{1}\times\Omega_{0}$. Thus $S$ splits into a disjoint union $S=S_{r}\cup S_{0,1}\cup S_{1,0}$ where $S_{r}$ stands for all regular relations while $S_{i,1-i}:=\{s\in S\mid\subseteq\Omega_{i}\times\Omega_{1-i}\}$.
Define $n_{s}:=\frac{1}{2}p_{\{s,s^{t}\}}^{1_{\Omega}}$, for each $s\in S_{i,1-i}$. Then $|s|=n_{s}|\Omega_{i}|$, and it follows from $\sum_{s\in S_{i,1-i}}|s|=|\Omega_{i}|\cdot|\Omega_{1-i}|$ that $\sum_{s\in S_{i,1-i}}n_{s}=|\Omega|/2$. Since all summands in the latter sum are nonnegative integers, we conclude $|S_{i,1-i}|\leq|\Omega|/2$ for $i=0,1$.
@@ 3.1. Nonregular thin Jordan schemes
$\mathcal{J}(\mathcal{C}):=\left\{\left[\begin{smallmatrix}A&B\\[2.84526pt] C&A\end{smallmatrix}\right]\;\middle|\;A,B,C\in\mathcal{C}\right\}.$
@@
Let $\mathfrak{J}=(\Omega,\mathcal{C})$ be a nonregular thin Jordan scheme and let $\mathcal{A}$ be its adjacency algebra. Let ${\mathcal{S}}=\{A_{i}\}_{i=1}^{r}$ be the standard basis of $\mathcal{A}$. By [14, Proposition 2.2], each element of $\mathcal{S}$} is of one of the following shapes:
@@
$\forall_{C\in\mathcal{H}}\,\forall_{A\in\mathcal{G}_{1}}\,\exists_{D\in\mathcal{K}}\,:\,\left[\begin{smallmatrix}A&O\\[2.84526pt] O&\varphi(A)\end{smallmatrix}\right]=\left[\begin{smallmatrix}CD&O\\[2.84526pt] O&DC\end{smallmatrix}\right].$ (2)
@@
$\forall_{C\in\mathcal{H}}\,\forall_{A\in\mathcal{G}_{1}}\,:\,\varphi(g)=C^{-1}AC.$ (3)
@@
Let $P$ be a permutation matrix and let $X\in{\mathcal{A}}$ be arbitrary. Then conjugation $X\mapsto PXP^{-1}$ preserves each of the operations $\cdot$, $\tau$, $\star$, $\circ$ on $\mathcal{A}$}. In other words, the resulting algebra $\mathcal{A}^{\prime}=P\mathcal{A}P^{-1}$ is a homogeneous coherent $J$-algebra and their corresponding Jordan schemes are combinatorially isomorphic.

### v3

@@ 1. Introduction
It is well known that all finite groups form a special subclass of association schemes, called thin schemes [23]. Such schemes may be characterized as those that have a maximum rank-to-order ratio among all association schemes. In this paper, we classify all Jordan schemes that attain maximum rank-to-order ratio. We further show that regular Jordan schemes with this property are in a one-to-one correspondence with a special class of Moufang loops. We direct the reader to [4] for detailed information about various types of loops, e.g. Moufang loops, Bol loops, etc.
The paper is organized as follows. Section 2 contains all necessary definitions and basic facts about coherent Jordan configurations and Jordan schemes. Section 3 is the most central part of our paper. There we show that the rank-to-order ratio of a Jordan scheme is bounded by $3/2$ with equality occurring if and only if the scheme is thin and nonregular. All such Jordan schemes are classified in Subsection 3.1. Subsection 3.2 treats thin regular Jordan schemes. There we establish a one-to-one correspondence between such schemes and ring-alternative Moufang loops (RA-loops). Finally, in Section 4 we show that Jordan schemes corresponding to RA-loops are autonomous, that is to say, they are not algebraic fusions of coherent configurations.
Thus we describe in Section 4 the first infinite family of autonomous Jordan schemes. Previously, the only known autonomous Jordan scheme was the one on $15$ points described in [18] (see also [14, Section 4]) related to one of the two Siamese $\mathrm{STS}(15)$ constructed in [16, Section 4], see also [17]. The fact that this Jordan scheme is autonomous was established using the COCO2P package [15] in GAP [8] and Ziv-Av’s catalog of small coherent configurations [24].
@@ 2.1. Binary relations and matrices
Let $\Omega$ be a nonempty finite set. A binary relation on $\Omega$ is an arbitrary subset $s\subseteq\Omega\times\Omega$. We refer to the set ${\omega}s:=\{\beta\in\Omega\mid(\omega,\beta)\in s\}$ as the $s$-neighborhood of $\omega\in\Omega$. A relation is called regular if $|\omega s|$ does not depend on $\omega$. In such case, we refer to the constant $n_{s}:=|\omega s|$ as the valency of $s$. If $s$ is a regular relation of valency one, then ${\omega}s$ contains a unique element which we denote as $s(\omega)$, i.e. ${\omega}s=\{s(\omega)\}$. Thus $s$ is a function on $\Omega$. All functions in this paper, including permutations from ${\sf Sym}(\Omega)$, will be considered as binary relations.
Every relation $s$ admits a transposed relation $s^{t}=\{(\alpha,\beta)\in\Omega^{2}\mid(\beta,\alpha)\in s\}$. Throughout, we will denote the unary operator that transposes relations by $\tau$, that is, $\tau:s\mapsto s^{t}$. Following [7] we call a relation $s\subseteq\Omega\times\Omega$ thin if $|\omega s|\leq 1\geq|\omega s^{t}|$ holds for each $\omega\in\Omega$.
@@
The algebra of all $\Omega\times\Omega$ matrices over a field $\mathbb{F}$ will be denoted as $M_{\Omega}(\mathbb{F})$. It is always assumed that $\mathsf{char}({\mathbb{F}})\neq 2$. The standard product of two matrices $A,B\in M_{\Omega}(\mathbb{F})$ will be written as $A\cdot B$ or $AB$, while their Jordan product $\frac{1}{2}(AB+BA)$ will be denoted by $A\star B$. The Schur-Hadamard (or entry-wise) product will be denoted as $A\circ B$. The identity matrix and all-ones matrix will be denoted as $I_{\Omega}$ and $J_{\Omega}$, respectively. The linear span of any set of matrices ${\mathcal{S}}\subseteq M_{\Omega}({\mathbb{F}})$ will be denoted by $\mathbb{F}[\mathcal{S}]$. In the case where $\mathcal{S}$ is given as a family $\{A_{i}\}_{i\in I}$, we also write $\langle A_{i}\rangle_{i\in I}$ for $\mathbb{F}[\mathcal{S}]$.
@@ 2.2. Rainbow algebras
Following [11] we call a pair $\mathfrak{S}=(\Omega,S)$ a rainbow if it satisfies these conditions. The elements of $S$ are called basic relations of the rainbow. The cardinalities $|\Omega|$ and $|S|$ are called the order and rank of the rainbow, respectively. Observe that a rainbow has rank one if and only if $|\Omega|=1$. For $|\Omega|\geq 2$ one has $|S|\geq 2$, where equality is attained if and only if $S=\{1_{\Omega},\,\Omega^{2}\setminus 1_{\Omega}\}$. Such rainbows are called trivial. The relation matrix of a rainbow ${\mathfrak{S}}=(\Omega,S)$ is the $\Omega\times\Omega$ matrix $A({\mathfrak{S}})$ defined via $A({\mathfrak{S}})_{\alpha,\beta}=S(\alpha,\beta)$ where $S(\alpha,\beta)$ is the unique basic relation containing $(\alpha,\beta)\in\Omega^{2}$.
@@ 2.3. Coherent Jordan algebras
Recall that a homogeneous coherent configuration is called an association scheme (see [23]). Following this pattern, we will call a homogeneous coherent J-configuration a Jordan scheme. While it is well known that association schemes are always regular, Jordan schemes do not possess this property in general. It was shown in [18] that a symmetric Jordan scheme is always regular. Regarding nonsymmetric Jordan schemes, they are not necessarily regular, but have a very special structure — see [18, Proposition 2.2]. Examples of nonregular Jordan schemes also appear in Subsection 3.1 of this paper.
@@ 2.5. Algebraic fusions of J-configurations
    For any $\Phi\leq\mathsf{JAut}({\mathfrak{S}})$, $\bigl(\Omega,\{\Phi(s)\mid s\in S\}\bigr)$ is a Jordan configuration.
@@
Proof of a: One readily checks that for any pair $r,s\in S$,
$s=r^{t}\iff(\underline{s}\star\underline{r})\circ I_{\Omega}\neq O_{\Omega}\mbox{ (the $\Omega\times\Omega$ zero matrix}).$
@@
$(\underline{s}\star\underline{r})\circ I_{\Omega}\neq O_{\Omega}\iff(\underline{\phi(s)}\star\underline{\phi(r)})\circ I_{\Omega}\neq O_{\Omega},$
@@
Proof of b: It follows from part (a) that
@@
Proof of c: Denote ${\mathcal{A}}:=\mathbb{C}[S]$ and extend each $\phi\in\Phi$ on $\mathcal{A}$} linearly via $\phi(\underline{s}):=\underline{\phi(s)}$. We will use the same notation for the induced linear operator on $\mathcal{A}$}. Then
@@ 2.6. Constructing Jordan configurations from coherent ones
If ${\mathfrak{S}}=(\Omega,S)$ is a coherent configuration, then it is also a Jordan configuration. Recall that a permutation $\phi$ of $S$ is called an algebraic automorphism (anti-automorphism) of $\mathfrak{S}$} if its linear extension to $\mathcal{A}:=\langle\underline{s}\rangle_{s\in S}$ is an automorphism (anti-automorphism) of $(\mathcal{A},\cdot)$. The group $\mathsf{AAut}({\mathfrak{S}})$ of all algebraic automorphisms of $\mathfrak{S}$} is contained in $\mathsf{JAut}({\mathfrak{S}})$. Also, $\tau\in\mathsf{JAut}({\mathfrak{S}})$. Every anti-automorphism of $\mathfrak{S}$} is a composition of $\tau$ and some $\phi\in\mathsf{AAut}({\mathfrak{S}})$. All automorphisms and anti-automorphisms of $\mathfrak{S}$} form a subgroup of $\mathsf{JAut}({\mathfrak{S}})$ which will be denoted by $\mathsf{TAut}({\mathfrak{S}})$. If $\mathfrak{S}$} is commutative, then $\mathsf{AAut}({\mathfrak{S}})=\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$. If however $\mathfrak{S}$} is not commutative, then $\tau\not\in\mathsf{AAut}({\mathfrak{S}})$ and $\mathsf{TAut}({\mathfrak{S}})\!:\!\mathsf{AAut}({\mathfrak{S}})]=2$ whence $\mathsf{TAut}({\mathfrak{S}})=\mathsf{AAut}({\mathfrak{S}})\times\langle\tau\rangle$. In general, the inclusion $\mathsf{TAut}({\mathfrak{S}})\leq\mathsf{JAut}({\mathfrak{S}})$ may be proper. The smallest examples of this are $\mathrm{CC}(8,31)$ and $\mathrm{CC}(8,32)$ of order $8$. (Enumeration of coherent configurations here is taken from Ziv-Av’s catalog [24] of small coherent configurations.)
If ${\mathfrak{S}}=(\Omega,S)$ is a thin association scheme, then $S$ is a regular subgroup of ${\sf Sym}(\Omega)$. In this case a permutation $\phi\in{\sf Sym}(S)$ belongs to $\mathsf{JAut}({\mathfrak{S}})$ if and only if it satisfies $\phi(st)\in\{\phi(s)\phi(t),\phi(t)\phi(s)\}$ for all $s,t\in S$. It was shown in [19] that any such map is either an automorphism or anti-automorphism. Thus for thin association schemes one always has $\mathsf{TAut}({\mathfrak{S}})=\mathsf{JAut}({\mathfrak{S}})$.
@@ 3. Thin Jordan schemes
By [18, Proposition 2.2] there exists a partition $\Omega=\Omega_{1}\cup\Omega_{0}$ with $|\Omega_{1}|=|\Omega_{0}|$ such that every regular relation is contained in $\Omega_{0}^{2}\cup\Omega_{1}^{2}$ and each nonregular one is contained in either $\Omega_{0}\times\Omega_{1}$ or $\Omega_{1}\times\Omega_{0}$. Thus $S$ splits into a disjoint union $S=S_{r}\cup S_{0,1}\cup S_{1,0}$ where $S_{r}$ stands for all regular relations while $S_{i,1-i}:=\{s\in S\mid\subseteq\Omega_{i}\times\Omega_{1-i}\}$.
Define $n_{s}:=\frac{1}{2}p_{\{s,s^{t}\}}^{1_{\Omega}}$ for each $s\in S_{i,1-i}$. Then $|s|=n_{s}|\Omega_{i}|$, and it follows from $\sum_{s\in S_{i,1-i}}|s|=|\Omega_{i}|\cdot|\Omega_{1-i}|$ that $\sum_{s\in S_{i,1-i}}n_{s}=|\Omega|/2$. Since all summands in the latter sum are nonnegative integers, we conclude $|S_{i,1-i}|\leq|\Omega|/2$ for $i=0,1$.
@@ 3.1. Nonregular thin Jordan schemes
$\mathcal{J}(\mathcal{C}):=\left\{\left[\begin{smallmatrix}A&B\\[5.69054pt] C&A\end{smallmatrix}\right]\;\middle|\;A,B,C\in\mathcal{C}\right\}.$
@@
Let $\mathfrak{J}=(\Omega,\mathcal{C})$ be a nonregular thin Jordan scheme and let $\mathcal{A}$ be its adjacency algebra. Let ${\mathcal{S}}=\{A_{i}\}_{i=1}^{r}$ be the standard basis of $\mathcal{A}$. By [18, Proposition 2.2], each element of $\mathcal{S}$} is of one of the following shapes:
@@
$\forall_{C\in\mathcal{H}}\,\forall_{A\in\mathcal{G}_{1}}\,\exists_{D\in\mathcal{K}}\,:\,\left[\begin{smallmatrix}A&O\\[5.69054pt] O&\varphi(A)\end{smallmatrix}\right]=\left[\begin{smallmatrix}CD&O\\[5.69054pt] O&DC\end{smallmatrix}\right].$ (2)
@@
$\forall_{C\in\mathcal{H}}\,\forall_{A\in\mathcal{G}_{1}}\,:\,\varphi(A)=C^{-1}AC.$ (3)
@@
Let $P$ be a permutation matrix and let $X\in{\mathcal{A}}$ be arbitrary. Then conjugation $X\mapsto PXP^{-1}$ preserves each of the operations $\cdot$, $\tau$, $\star$, $\circ$ on $\mathcal{A}$}. In other words, the resulting algebra $\mathcal{A}^{\prime}=P\mathcal{A}P^{-1}$ is a homogeneous coherent $J$-algebra, and the Jordan schemes corresponding to $\mathcal{A}$ and $\mathcal{A^{\prime}}$ are combinatorially isomorphic.
@@
Let us collect our findings in one place:
Theorem 3.5.
  1. (a)
    For every finite abelian group $G$, $\mathcal{J}(G)$ is the adjacency algebra of a non-regular thin Jordan scheme.
  2. (b)
    For every non-regular thin Jordan scheme $\mathfrak{S}$ with adjacency algebra $\mathcal{A}$}, there exists a finite abelian group $G$ such that ${\mathcal{A}}\cong\mathcal{J}(G)$.
  3. (c)
    For finite abelian groups $G$ and $H$, one has $\mathcal{J}(G)\cong\mathcal{J}(H)$ if and only if $G\cong H$.
Proof.
Parts (a) and (b) follow from Propositions 3.3 and 3.4, respectively. For part (c), clearly, if $G\cong H$ then $\mathcal{J}(G)\cong\mathcal{J}(H)$. For the reverse direction we note that $G$ can be reconstructed from $\mathcal{J}(G)$ as follows: observe that the WL-closure $\mathcal{W}$ of $\mathcal{J}(G)$ is equal to
$\mathcal{W}=\left\{\left[\begin{smallmatrix}A&B\\[5.69054pt] C&D\end{smallmatrix}\right]\;\middle|\;A,B,C,D\in{\mathcal{A}}(G,G)\right\}.$
In other words it is a coherent algebra with two fibres where each fibre constituent coincides with ${\mathcal{A}}(G,G)$. ∎